Given the equation, 9y + 36 - 4x = 0, find the slope and the y-intercept.
Answer by tutorcecilia(2144)

9y + 36 - 4x = 0, find the slope and the y-intercept
9y + 36 -36 - 4x + 4x= 0 -36 + 4x
9y= -36 + 4x
9y/9 = -36/9 + 4x/9
y= -4 + 4/9x
y = 4/9x - 4
Slope - m = 4/9
Y-intercept = b = -4